if a rifle fires a bullet of mass 0.01kg with a velocity of 300m s-1 and the rifle has a mass of 60kg. Determine the velocity of the rifle if it recoils or kick back.
1
Expert's answer
2016-10-20T11:19:09-0400
p1 = p2
m1*v1 = m2*v2
V2 = m1*v1/m2
V2 = 0.01*300/60 = 0.05 m/s